# A flexible and general semi-supervised approach to multiple hypothesis testing

### v1

@@ 2.1 Extension to side information
In proteomics, so-called post-processors such as PeptideProphet [28] and Percolator [26] have popularized the use of side-information to improve the number of detected peptides. These tools use machine learning procedures, leveraging features associated with the PSMs, such as the charge state associated with the spectrum, and the matched peptide’s length, to better distinguish correct PSMs, and hence to deliver more discoveries. These tools have no theoretical guarantees and having recently pointed out scenarious where such post-processors can fail to control the FDR [18], we will not analyze these tools in this paper.
The only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner, that we are aware of, is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès [43]. AdaKO is a specialization of the p-value based procedure AdaPT [33] that we discuss in the next section. By ‘intrinsic’, we mean that while other p-value based approaches may be modified to be used in the competition setting, they are significantly less powerful and do not really qualify in this sense, as demonstrated in [43]. Adaptive Knockoffs works by iteratively pruning $\mathcal{S}$, the candidate set of hypotheses, selecting each time the hypothesis it deems most likely to be a true null. To do that, it fits a classification algorithm using a strict set of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, the labels $L_{i}$ of the hypotheses that were pruned up to that point ($i\in[p]\setminus\mathcal{S}$), as well as #$\#\{L_{i}=1:i\in S\}$ and #$\#\{L_{i}=-1:i\in S\}$. In their paper, Ren and Candès investigate this prunning strategy using different types of “filters”: a logistic regression (LR), a generative additive model (GAM) which is the default filter in their R package, a random forest (RF) and a two-group probabilistic model which is fitted using the expectation-maximization algorithm (EM) [13]. The pruning process terminates when the estimated FDR is less than or equal to $\alpha$:
@@
The remaining positively labelled hypotheses in $\mathcal{S}$ are then reported. It was shown that given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
@@
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ uniform random variables independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
@@ 3 Background: p-value based multiple testing
Multiple testing with FDR control was popularized by Benjamini and Hochberg when they developed the first testing procedure of this kind, which we refer to as the “BH procedure” [2]. Since then, many p-value based methods that control the FDR have been developed to improve statistical power.
One such strategy for improving on BH is to estimate the fraction of true null hypotheses, denoted as $\pi_{0}$. Indeed, the BH procedure controls the FDR conservatively at level $\pi_{0}\cdot\alpha$ instead of $\alpha$. Hence, in cases when $\pi_{0}\ll 1$, the BH procedure loses a considerable amount of power. Methods like those by Storey et al. [45] and Benjamini et al. [3] estimate $\pi_{0}$, and essentially ‘plug-in’ the estimate $\hat{\pi}_{0}$, by applying the BH procedure at a threshold of $\alpha/\hat{\pi}_{0}$ instead of $\alpha$.
Instead of the FDR, an alternative to BH that controls the FDP, is Guo and Romano’s stepdown approach [20], which we refer to as GR-SD. In fact, this stepdown approach inspired the development of FDP-SD in the competition setting. GR-SD works by first sorting the p-values in ascending order, i.e. $p_{1}\leq p_{2}\leq...\leq p_{m}$. Then it constructs a sequence of thresholds, $\delta_{i}$ for each p-value $p_{i}$ and searches for the largest $k$ such that $p_{1}\leq\delta_{1},\dots,p_{k}\leq\delta_{k}$ mutually holds, and then reports the $k$ smallest p-values. We provide the pseudocode for this method in Algorithm S3.
@@ 3.1 Extension to side information
Other methods utilize side-information in one of two ways: (a) they provide an improved ordering of the p-values based on how confident we believe each hypothesis is a false null, or (b) by assigning weights to the p-values, placing higher weights on hypotheses that we believe to be false nulls. However, the shortcoming of these side-information approaches is that the majority of them are not data-driven, or ‘adaptive’, that is, the ordering or weighting is fixed before the actual testing procedure takes place.
@@
Table 1: A table summarizing some popular and recent p-value based multiple testing procedures that use side information in terms of its type-1 error control and the computational speed.
@@
More recently, several methods have been developed that take an adaptive approach instead. Such methods include Adaptive P-value Thresholding (AdaPT) and its derivatives [33, 8], Z-value Adaptive Procedures (ZAP) [35] and AdaFDR [50]. Table 1 provides a qualitative summary of each method in terms of its type-1 error control, its error control guarantees, and its computational speed. On balance, none of these mentioned tools enjoy both finite type-1 error control while also being fast, and as we will see, some of these methods have variable power when considering a range of simulated and real datasets. We next provide some background on each of those methods.
@@
AdaFDR, splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds. For each fold, it fits a mixture model to optimally rescore the hypotheses, which is then applied to the other fold where a list of discoveries is obtained. The two list of discoveries are joined together and then reported. AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus a generalized linear model to capture a monotonic relationship between the side information and the signals.
@@
AdaPT iteratively prunes a candidate set of hypotheses $\mathcal{S}$ similar to Adaptive Knockoffs. At each iteration, AdaPT fits a two group mixture model on a strict set of information: the p-values outside $\mathcal{S}$, the masked p-values in $\mathcal{S}$ given by $\{p_{i}\wedge(1-p_{i}):i\in\mathcal{S}\}$, the side information $\mathbf{x}$, as well as #$\#\{p_{i}>1/2:i\in\mathcal{S}\}$ and #$\#\{p_{i}<1/2:i\in\mathcal{S}\}$. It then removes the hypothesis from $\mathcal{S}$ that is deemed most likely to be a true null based on the fitted model. To fit the mixture model, the EM algorithm is used. The M-step of the EM algorithm results in a regression problem which allows the user to flexibly choose whichever regression method they like. They considered a generative additive model (GAM), a generalized linear model (GLM) and a generalized linear model with $L_{1}$ regularization (GLMnet). If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure when the estimated FDR, like in Equation (4), is $\leq\alpha$. The rationale is that for true null hypotheses, the p-values are usually uniformly distributed, so #$\#\{L_{i}=-1:i\in\mathcal{S}\}$ estimates the number of true null hypotheses with positive labels, #$\#\{L_{i}=1:i\in\mathcal{S}\}$. It then reports the remaining positively labelled hypotheses in $\mathcal{S}$ [^3] AdaPT and AdaPTg do not use labels explicitly – we introduce those here for notational convenience.
@@
AdaPTg generalizes AdaPT by constructing asymmetric regions to define the positive and negative labels. For example, the p-values $p_{i}\in\left(0.3,0.9\right)$ can be identified with a label $L_{i}=-1$ and the p-values $p_{i}<0.3$ can be identified with a label $L_{i}=1$. One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses when there is a concentration of p-values towards 1. Subsequently, the estimated FDR needs to be adjusted to account for the fact that we expect twice as many true null p-values with negative labels than positive labels. Moreover, the masked p-values are also no longer $p_{i}\wedge(1-p_{i})$ and also need adjustment. In essence, the rest of the procedure remains the same. Lastly, AdaPT-GMMg replaces the two group mixture model in AdaPT/AdaPTg for a Gaussian mixture model (GMM). Similar to AdaPT/AdaPTg, it also allows the user decide from a collection of regression methods to fit the model at the M-step of the EM algorithm.
@@
Lastly, ZAP-asymp directly fits a beta mixture model to the data. ZAP-asymp’s direct application of the model means that it is only able to offer asymptotic FDR control. Importantly, ZAP-asymp can operate directly on the test statistic instead of just the corresponding p-values to rerank the hypotheses. This is particularly advantageous for two-sided tests because the test statistic is mapped to its p-value in a non-bijective manner, losing information regarding the sign of the original test statistic along the way. Note that AdaPT-GMMg can also operate on test statistics, but in this paper, we focus on the p-value information.
@@ 4 RESET (REScoring via Estimating and Training)
RESET is designed to control the FDR or the FDP in either the competition or p-value based setting. We next describe RESET’s main steps, with further details given in Algorithm S4. In Section S.3, we prove that RESET controls the FDR/FDP in the finite-sample setting. For simplicity, we use the target-decoy terminology, i.e., we refer to a hypothesis with $L=1$ as a ‘target win’ or a target for short, and $L=-1$ as a ‘decoy win’ or a decoy for short.
@@ 4.1 RESET outline
    Intrinsically, RESET’s input consists of the labels, winning scores and side information $(L,W,\mathbf{x})$. Hence, in the p-value setting, each $p_{i}$ is first converted to a pair $(L_{i},W_{i})$ in the following way:
@@
    where $0<a\leq 1/2\leq b_{1}<b_{2}\leq 1$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF. Hypotheses with p-values outside of $[0,a)\cup(b_{1},b_{2}]$ are thrown out and the default is $a=b_{1}=1/2$ and $b_{2}=1$.
@@
    RESET independently and randomly assigns each winning decoy as a training decoy with probability $s$ (we used $s=1/2$ throughout). The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo targets is the set of all estimating decoys and target wins. The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
@@
    Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores and side information $(\tilde{L},W,\mathbf{x})$. Note that this is a semi-supervised task as the positive pseudo labels will typically contain a mixture of true and false nulls, while the negative pseudo labels will indicate true null hypotheses.
    The output of this step is a rescoring of the training decoys and pseudo targets, denoted as $\tilde{W}$. Ideally, $\tilde{W}$ scores many of the false nulls among the pseudo targets higher than they were scored originally. We provide a specific framework that we developed for this crucial step in Section 4.3.
@@
    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo targets are revealed.
@@
    RESET then determines its list of discoveries, by applying to the pseudo-targets using the original labels $L$ and the new scores $\tilde{W}$ either
@@
    The two algorithms, SSS+ and FDP-SD, require an additional parameter $c$. This parameter is used to define the expected ratio of the number of null targets to decoys, $\frac{c}{1-c}$. Typically, this is set to $c=1/2$ for an expected target-decoy ratio of 1, however, RESET uses approximately half of the decoys for training purposes. Since those decoys are subsequently thrown out, this parameter needs to be adjusted. Indeed, assuming that the labels $L$ of the true nulls are i.i.d uniform ±$\pm 1$ RVs, and $s=1/2$, we set the parameter $c=2/3$, so that the expected target-decoy ratio is 2. More generally, if we know $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$. For example in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$. This choice is made rigorous in Lemma S1.
We note that the above application of $\Phi^{-1}$ in Equation (5) is not strictly necessary for type-1 error rate control. We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$ because we found that doing so makes the semi-supervised learning task of Step 3 to be generally less variable and more powerful (data not shown). The idea of mirrored p-value was established by Lei and Fithian [33] and later generalized by [8] as discussed in the previous section.
Crucially in Step 4, the training decoys are thrown out. This is important since during the training phase of Step 3, the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo targets. Hence, the training decoys are no longer suitable for estimating the true nulls among those pseudo targets, because the rescoring will tend to rank them lower than those true nulls.

### v2

@@ 2.1 Extension to side information
In proteomics, so-called post-processors such as PeptideProphet [28] and Percolator [26] have popularized the use of side information to increase the number of detected peptides. These tools use machine learning procedures, leveraging features associated with the PSMs, such as the charge state associated with the spectrum and the matched peptide’s length, to better distinguish correct from incorrect PSMs, and hence to deliver more discoveries. These tools have no theoretical guarantees and, having recently pointed out scenarios where such post-processors can fail to control the FDR [18], we will not analyze these tools in this paper.
To our knowledge, the only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès [44]. AdaKO is a specialization of the p-value based procedure AdaPT [33] that we discuss in the next section. By ‘intrinsic’, we mean that while other p-value based approaches may be modified to be used in the competition setting, they are significantly less powerful and do not really qualify in this sense, as demonstrated in [44]. Adaptive Knockoffs works by iteratively pruning $\mathcal{S}$, the candidate set of hypotheses, selecting each time the hypothesis it deems most likely to be a true null. To do that, it fits a classification algorithm using a strict set of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, and the labels $L_{i}$ of the hypotheses that were pruned up to that point ($i\in[p]\setminus\mathcal{S}$), as well as #$\#\{L_{i}=1:i\in S\}$ and #$\#\{L_{i}=-1:i\in S\}$. In their paper, Ren and Candès investigate this pruning strategy using different types of ‘filters’: a logistic regression (LR), a generalized additive model (GAM) which is the default filter in their R package, a random forest (RF), and a two-group probabilistic model which is fitted using the expectation-maximization algorithm (EM) [13]. The pruning process terminates when the estimated FDR is less than or equal to $\alpha$:
@@
The remaining positively labeled hypotheses in $\mathcal{S}$ are then reported. It was shown that, given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
@@
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d. ±$\pm 1$ uniform random variables independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
@@ 3 Background: p-value based multiple testing
Multiple testing with FDR control was popularized by Benjamini and Hochberg when they developed the first testing procedure of this kind, which we refer to as the ‘BH procedure’ [2]. Since then, many p-value based methods that control the FDR have been developed to improve statistical power.
One such strategy for improving on BH involves estimating the fraction of true null hypotheses, denoted as $\pi_{0}$. Indeed, the BH procedure controls the FDR conservatively at level $\pi_{0}\cdot\alpha$ instead of $\alpha$. Hence, in cases when $\pi_{0}\ll 1$, the BH procedure loses a considerable amount of power. Methods like those by Storey et al. [46] and Benjamini et al. [3] estimate $\pi_{0}$, and essentially ‘plug-in’ the estimate $\hat{\pi}_{0}$, by applying the BH procedure at a threshold of $\alpha/\hat{\pi}_{0}$ instead of $\alpha$.
Instead of the FDR, Guo and Romano’s stepdown approach [20], which we refer to as GR-SD, controls the FDP. In fact, this stepdown approach inspired the development of FDP-SD in the competition setting. GR-SD works by first sorting the p-values in ascending order, i.e., $p_{1}\leq p_{2}\leq...\leq p_{m}$. Then it constructs a sequence of thresholds, $\delta_{i}$ for each p-value $p_{i}$ and searches for the largest $k$ such that $p_{1}\leq\delta_{1},\dots,p_{k}\leq\delta_{k}$ mutually holds, and then reports the corresponding $k$ smallest p-values. We provide the pseudocode for this method in Algorithm S3.
@@ 3.1 Extension to side information
Subsequent approaches were developed to take advantage of scenarios with available side information. Roughly speaking, these methods use side information to weight or rescore the hypotheses, placing higher weights or scores on hypotheses that are believed to be false nulls. Such procedures can be either ‘predetermined’, that is, the weighting or scoring is determined before the actual testing procedure takes place, or ‘adaptive’, in which the weighting or scoring is selected according to the data at hand and therefore often leads to greater statistical power.
@@
Table 1: Summary of some popular and recent p-value based multiple testing procedures that use side information, indicating each method’s type-1 error control and speed.
@@
Such adaptive methods include Adaptive P-value Thresholding (AdaPT) and its derivatives [33, 8], Z-value Adaptive Procedures (ZAP) [35] and AdaFDR [51]. Table 1 provides a qualitative summary of each method in terms of its type-1 error control, its error control guarantees, and its computational speed. Notably, none of these mentioned tools offers a fast finite type-1 error control, and as we will see, some of these methods have variable power when considering a range of simulated and real datasets. We next provide some background on each of those methods.
@@
AdaFDR splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds. For each fold, the method fits a mixture model to optimally rescore the hypotheses, and the model is then applied to the other fold where a list of discoveries is obtained. The two list of discoveries are joined together and then reported. AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus an additional component that captures the monotonic relationship between the side information and the signals.
@@
AdaPT iteratively prunes a candidate set of hypotheses $\mathcal{S}$, similar to Adaptive Knockoffs, which was derived from it. At each iteration, AdaPT fits a two-group mixture model on a strict set of information: the p-values outside $\mathcal{S}$, the masked p-values in $\mathcal{S}$ given by $\{p_{i}\wedge(1-p_{i}):i\in\mathcal{S}\}$ where $1-p_{i}$ is referred to as the mirrored p-value, the side information $\mathbf{x}$, as well as #$\#\{p_{i}>1/2:i\in\mathcal{S}\}$ and #$\#\{p_{i}<1/2:i\in\mathcal{S}\}$. It then removes the hypothesis from $\mathcal{S}$ that is deemed most likely to be a true null based on the fitted model. To fit the mixture model, the EM algorithm is used. The M-step of the EM algorithm results in a regression problem which allows the user to flexibly use the regression method of their choice. The authors considered a generalized additive model (GAM), a generalized linear model (GLM), and a generalized linear model with $L_{1}$ regularization (GLMnet). If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure when the estimated FDR, as in Equation (4), is $\leq\alpha$. The rationale is that for true null hypotheses, the p-values are usually uniformly distributed, so #$\#\{L_{i}=-1:i\in\mathcal{S}\}$ estimates the number of true null hypotheses with positive labels, #$\#\{L_{i}=1:i\in\mathcal{S}\}$. It then reports the remaining positively labeled hypotheses in $\mathcal{S}$ [^3] AdaPT and AdaPTg do not use labels explicitly — we introduce those here for notational convenience.
@@
AdaPTg generalizes AdaPT by constructing asymmetric regions to define the positive and negative labels. For example, the p-values $p_{i}\in\left(0.3,0.9\right)$ can be identified with a label $L_{i}=-1$ and the p-values $p_{i}<0.3$ can be identified with a label $L_{i}=1$. One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses in cases where their p-values are not uniformly distributed and there is a concentration of true null p-values near 1. Subsequently, the estimated FDR needs to be adjusted to account for the fact that we expect twice as many true null p-values with negative labels than positive labels. Moreover, the masked p-values are also no longer $p_{i}\wedge(1-p_{i})$, since $1-p_{i}$ no longer ‘correctly mirrors’ the values in $\left(0.3,0.9\right)$ onto $(0,0.3)$ and therefore needs adjustment. In essence, the rest of the procedure remains the same. Lastly, AdaPT-GMMg replaces the two group mixture model in AdaPT/AdaPTg for a Gaussian mixture model (GMM). Similar to AdaPT/AdaPTg, it also allows the user to choose from a collection of regression methods to fit the model at the M-step of the EM algorithm.
@@
Lastly, ZAP-asymp directly fits a beta mixture model to the data but is only able to offer asymptotic FDR control. Importantly, ZAP-asymp can operate directly on the test statistic instead of just the corresponding p-values to rerank the hypotheses. This is particularly advantageous for two-sided tests because the test statistic is mapped to its p-value in a non-bijective manner, losing information regarding the sign of the original test statistic along the way. Note that AdaPT-GMMg can also operate on test statistics, but in this paper, we focus on the p-value information.
@@ 4 RESET (REScoring via Estimating and Training)
RESET is designed to control the FDR or the FDP in either the competition or p-value based setting. We next describe RESET’s main steps, with further details given in Algorithm S4. In Section S.3, we prove that RESET controls the FDR/FDP in the finite-sample setting. For simplicity, we use the target-decoy terminology, i.e., we refer to a hypothesis with $L_{i}=1$ as a ‘target win’ or a target for short, and $L_{i}=-1$ as a ‘decoy win’ or a decoy for short.
@@ 4.1 RESET outline
    Intrinsically, RESET’s input consists of the labels, winning scores, and side information $(L,W,\mathbf{x})$. Hence, in the p-value setting, each $p_{i}$ is first converted to a pair $(L_{i},W_{i})$ in the following way:
@@
    where $0<a\leq 1/2\leq b_{1}<b_{2}\leq 1$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF. Hypotheses with p-values outside of $[0,a)\cup(b_{1},b_{2}]$ are thrown out. The default is $a=b_{1}=1/2$ and $b_{2}=1$.
@@
    RESET independently and randomly assigns each winning decoy as a training decoy with probability $s$ (we used $s=1/2$ throughout). The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo-targets is the set of all estimating decoys and target wins. The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo-target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
@@
    Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores, and side information $(\tilde{L},W,\mathbf{x})$. Note that this is a semi-supervised task because the positive pseudo labels typically contain a mixture of true and false nulls, while the negative pseudo labels indicate true null hypotheses.
    The output of this step is a rescoring of the training decoys and pseudo-targets, denoted as $\tilde{W}$. Ideally, $\tilde{W}$ scores many of the false nulls among the pseudo-targets higher than they were scored originally. We provide a specific framework that we developed for this crucial step in Section 4.3.
@@
    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo-targets are revealed.
@@
    RESET then determines its list of discoveries by applying to the pseudo-targets using the original labels $L$ and the new scores $\tilde{W}$ either
@@
    The two algorithms, SSS+ and FDP-SD, require an additional parameter $c$. This parameter is used to define the expected ratio of the number of null targets to decoys, $\frac{c}{1-c}$. Typically, this is set to $c=1/2$ for an expected target-decoy ratio of 1; however, RESET uses approximately half of the decoys for training purposes (by default). Since those decoys are subsequently thrown out, this parameter needs to be adjusted. Indeed, assuming that the labels $L$ of the true nulls are i.i.d. uniform ±$\pm 1$ RVs, and $s=1/2$, we set the parameter $c=2/3$, so that the expected target-decoy ratio is 2. More generally, if we know that for a true null $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$. For example in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$. This choice is made rigorous in Lemma S1.
We note that the above application of $\Phi^{-1}$ in Equation (5) is not strictly necessary for type-1 error rate control. We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$ because we found that doing so makes the semi-supervised learning task of Step 3 to be generally less variable and more powerful (data not shown).
Crucially in Step 4, the training decoys are thrown out. This is important since during the training phase of Step 3, the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo-targets. Hence, true null targets will tend to score higher than the training decoys making the latter unsuitable for estimating the number of true nulls among the top scoring hypotheses.